OpenStudy (sidney18):

a(0.3−y)+1.1+2.4x= 0 (y−1.2)=−1.2(x−0.5) Consider the system of equations above, where a is a constant. For which value of a is there exactly one (x, y) solution where y = 1.3y? Round your answer to the nearest tenth.

OpenStudy (michele_laino):

we can rewrite the second equation, as follows: \[\begin{gathered} y - 1.2 = - 1.2x + 6 \hfill \\ \hfill \\ y = - 1.2x + 7.2 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

next I substitute such value of \(y\), into the first equation, namely I am applying the substitution method: \[0.3\left( {a - \left( { - 1.2x + 7.2} \right)} \right) + 1.1 + 2.4x = 0\] please simplify the left side

OpenStudy (michele_laino):

oops... sorry I have made a typo: \[a\left( {0.3 - \left( { - 1.2x + 7.2} \right)} \right) + 1.1 + 2.4x = 0\] please simplify

OpenStudy (michele_laino):

hint: if we apply the properties of the algebra of real numbers, namely the distributive property of multiplication, we get: \[1.2\left( {a + 2} \right)x = 6.9 - 0.3a\]

OpenStudy (michele_laino):

the condition: \(y=1.3y\) means \(1.3y-y=0\), so we get: \(0.3y=0\), whose solution is \(y=0\). Next I substitute \(y=0\) into the second equation: \((0-1.2)=-1.2(x-0.5)\) from which I get: \(-1.2=-1.2x+0.6\) and \(x=1.5\) finally I substitute the values \(x=1.5,y=0\) into the first equation, and I get: \[a\left( {0.3 - 0} \right) + 1.1 + \left( {2.4 \cdot 1.5} \right) = 0\] please solve for \(a\)

OpenStudy (sidney18):

a=0? because after you simplify you are left with a(5)=0 or at least that's what I got

OpenStudy (michele_laino):

I got this: \[0.3a + 4.7 = 0\] so, we have: \[a = \frac{{ - 4.7}}{{0.3}} = ...?\] please complete